Plot the point whose spherical coordinates are given. Then find the rectangular coordinates of the point. (a) (7,7/3, 7/6) 3(b) (3, п/2, 3/4) ы Зл 3 л 4 T 3 о 3 л T 2 у y Зл 3 л 4 4 (x, y, 2) =
Answer #1

Solution -  

a)   

(7, 3 6

Comparing it with (r,\theta ,\phi )

r=7 ,\theta =\frac{\pi }{3},\phi =\frac{\pi }{6}

   r is radius or distance of point from origin.\theta is the azimuthal angle in xy plane and  \phi is the polar angle.

So it is graph of point -

  TH J

x=rcos\theta sin\phi = 7cos(\frac{\pi }{3})sin(\frac{\pi }{6})

= 7(\frac{1}{2})(\frac{1}{2})=\frac{7}{4}

  y=rsin\theta sin\phi =7sin(\frac{\pi }{3})sin(\frac{\pi }{6})

  =7(\frac{\sqrt{3}}{2})(\frac{1}{2})=\frac{7\sqrt{3}}{4}

z=rcos\phi

=7 cos(\frac{\pi }{3})=7(\frac{1}{2})= \frac{7}{2}

So rectangular coordinates of point is (\frac{7}{4},\frac{7\sqrt{3}}{4},\frac{7}{2})

b)

(3,\frac{\pi }{2},\frac{3\pi }{4})

  r=3,\theta =\frac{\pi }{2},\phi =\frac{3\pi }{4}

So graph is following -

  1595125452535_image.png

x=rcos\theta sin\phi = 3cos(\frac{\pi }{2})sin(\frac{3\pi }{4})

=0

y=rsin\theta sin\phi =3sin(\frac{\pi }{2})sin(\frac{3\pi }{4})

=\frac{3\sqrt{2}}{2}

z=rcos\phi =7cos\left ( \frac{3\pi }{4} \right )=-\frac{3\sqrt{2}}{2}

So rectangular coordinates of point is (0,\frac{3\sqrt{2}}{2},-\frac{3\sqrt{2}}{4}) Answer